# LLLMs: A Data-Driven Survey of Evolving Research on Limitations of Large Language Models

### v2

@@ 3.4. Models and Prompting Evaluation
Similar to human annotators highlighting evidence in abstracts to indicate discussions of LLLMs, models are prompted to extract supporting evidence from the text. Each prompt instructs the model: ?Please respond in the following format, providing a rating and supporting evidence for the discussion of LLM limitations in each abstract. Do not include explanations, only cite the evidence found in the abstract.? Prompt 3, included as the most comprehensive, is available in Figure 14 of the supplementary material.
@@ 3.5. Clustering
To ensure meaningful clusters and minimize spurious outliers, we tune UMAP, HDBSCAN and BERTopic parameters separately for ACL and arXiv. We further apply a distance-based outlier reassignment strategy after observing that many outliers were not true noise but semantically close to existing clusters, suggesting misclassification by HDBSCAN. Full parameter settings and details of the reassignment procedure are provided in Section B of the supplementary materials.
@@ 4.1. LLM-based filtering evaluation
(1) ?Unlocking Adversarial Suffix Optimization Without Affirmative Phrases: Efficient Black-box Jailbreaking via LLM as Optimizer? (Jiang et al., 2024a),
@@ 4.2. LLM and LLLMs Trends Over Time
In both corpora, the (i) overall share of LLM-related papers has grown substantially since early 2023. This trend is particularly steep in ACL, where, by late 2024, over 75% of ACL papers are related to LLMs. This suggests a notable shift in NLP research, with LLMs becoming central to the field. In arXiv, growth is more moderate but consistent, hitting just above 30% of papers by the end of the same period. The lower proportion in arXiv might be due to different levels of engagement with LLM research across the categories in our study, as shown in Figure 17 in the supplementary material. In cs.CL, LLMs are widely discussed, reaching 80% by early 2025, similar to ACL, while in areas like cs.CV and cs.LG, their presence remains below 20%.
@@
ii Proportion of LLM limitation papers among all LLM papers.
@@
In the following sections, we refine this analysis and examine these emerging discussions in detail through topic clustering.
4.3. Clustering Results (HDBSCAN + BERTopic)
4.3.1. Topics identified within ACL and arXiv with HDBSCAN
@@
i ACL dataset
ii arXiv dataset
Figure 7. Wordclouds for LLM limitation topics identified by HDBSCAN+Bertopic, generated based on TF-IDF scores computed over concatenated evidence and keyphrases grouped by topic.
Key Insights HDBSCAN identifies 7 limitation topics in ACL and 15 in arXiv, with shared themes including Reasoning, Social Bias, Security Risks, and Hallucinations. ACL’s limitation research is dominated by Reasoning (36.4%), while arXiv presents a broader topical spread, including specialized areas like Healthcare, Code, and Quantization.
Figure 6 presents the distribution of LLM limitation topics in the ACL Anthology and arXiv datasets, as identified by the HDBSCAN + BERTopic clustering approach. Across both corpora, the model identified 7 topics for ACL and 15 for arXiv, in addition to a set of outliers (13.1% in ACL, 14.1% in arXiv).
We find that, according to the HDBSCAN approach, ACL’s limitation discussions are dominated by a single topic: Reasoning, which accounts for over one-third (36.4%) of all limitation-focused papers. In contrast, arXiv presents a more balanced topical spread, with the most frequent themes being Social Bias (16.8%), Security Risks (15.9%), and Reasoning (10.9%).
Several topics are shared across both corpora, including Hallucination, Social Bias, and Security Risks. However, arXiv also introduces a number of unique topics not observed in ACL, such as Multimodality, Healthcare Applications, Multilinguality, and Computational Cost.
Figure 7i and Figure 7ii show the dominant terms for each limitation cluster in the ACL and arXiv datasets; additionally, Table 7 and Table 8 present representative paper examples for some of the ACL and arXiv clusters, respectively. In ACL, the Reasoning cluster includes core cognitive tasks such as natural language understanding (NLU), inference, and logical analysis and reflects a broad range of reasoning types (e.g., temporal, causal, etc.). It spans multimodal tasks (example 1 in Table 7), visual question answering, and overlaps with multilinguality and benchmark design. In arXiv, Reasoning appears more distinct from other topics, though it remains closely linked to NLU, as shown by keywords like understanding and cognitive. Additionally, it includes prompt-related terms such as chain-of-thought and prompt, pointing to the role of prompt engineering in enabling complex reasoning. The content differences between ACL (wider topic) and arXiv (narrower topic) may also explain why the reasoning cluster is much more prevalent in ACL compared to arXiv, though this may also be a clustering artefact from HDBSCAN handling datasets differently.
@@
Several clusters focus on the correctness and reliability of model outputs. The Hallucination cluster addresses factual accuracy, including terms related to faithfulness, trust, and correctness (e.g., example 2). It also appears in multimodal contexts, with frequent references to hallucinations in image captioning and generation. The Security cluster highlights threats such as jailbreaks, adversarial prompts, and backdoors (e.g., example 3 shows adversarial attacks that evade detectors using prompt and synonym manipulation).
The Social Bias cluster focuses on fairness, representation, and demographic disparity, as evidenced by keywords such as fairness, representation, gender, racial, cultural, and demographic. It also includes terms tied to ethical and political considerations, including ethical, moral, political, and societal. In arXiv, the prominence of prompt, sensitivity, reliability, and variability suggests an emphasis on prompt-based bias mitigation (Xu et al., 2024b).
@@
Other clusters reflect technical constraints in model design and deployment. These include Long Context (ACL) and Context & Memory Limitations (arXiv), which address failures on long inputs and outputs, memory constraints, and, as a result, computational costs (example 6 in Table 7 and example 8 in Table 8). The Generalization cluster captures similar issues related to robustness and domain transfer (example 4).
Some clusters are specific to one dataset. In ACL, the Uncertainty cluster describes behavioral instability, including prompt sensitivity, calibration errors, and limited self-correction. In arXiv, additional clusters capture a wider range of specialized concerns: Multimodality, Conversational Limitations, Multilinguality, Code Generation, Healthcare Application, Computational Cost, Benchmark Contamination, Knowledge Editing, and Quantization. These clusters reflect both domain-specific challenges (e.g. Healthcare Application, which focuses on the use of LLMs in clinical settings and is primarily concerned with their black-box nature, see example 10) and implementation-level trade-offs. For example, Benchmark Contamination refers to test data leakage into training sets (example 12), Knowledge Editing focuses on updating model content without retraining (example 14), and Quantization and Computational Cost concern scaling and deployment efficiency (examples 13 and 11).
Table 7. Representative limitation-focused ACL papers, grouped by topic based on HDBSCAN clustering. Evidence is extracted from papers with a topic probability of 0.9 or higher. The numbers in parentheses in the ?Topic? column indicate the example ID referenced in the main text.
@@
Hallucination
?Challenges on hallucination and factual inconsistency continue to impede their [LLMs’] wider real-world adoption. […] However, challenges remain, particularly regarding… generating information not present in the evidence (hallucination).? (EACL 2024 (Lv et al., 2024))
Security
?The prevalence and strong capability of large language models (LLMs) present significant safety and ethical risks if exploited by malicious users. […] Experiments reveal that our attacks effectively compromise the performance of all detectors in the study with plausible generations […].? (TACL 2024 (Shi et al., 2024))
@@
?LLMs are generally trained on publicly available text and code and cannot be expected to directly generalize to domain-specific parsing tasks in a zero-shot setting […].? (EMNLP 2023 (Mekala et al., 2022))
Social Bias
@@
Uncertainty
[…] ?We find that methods aimed at improving usability, such as fine-tuning and chain-of-thought (CoT) prompting, can lead to miscalibration and unreliable natural language explanations.? (NAACL 2024 (Zhang et al., 2023c))
Table 8. Representative limitation-focused arXiv papers, grouped by topic based on HDBSCAN clustering. Only arXiv-specific clusters are shown; topics shared with ACL are excluded to avoid redundancy. Examples are selected based on a topic probability of 0.9 or higher. The numbers in parentheses in the ?Topic? column indicate the example ID referenced in the main text.
Topic Evidence
(8) Context &
Memory
?Experiments on CoIN demonstrate that current powerful MLLMs still suffer catastrophic forgetting, and the failure in intention alignment assumes the main responsibility, instead of the knowledge forgetting.? (arXiv, March 2024 (Chen et al., 2024b))
(9) Conversational
Limitations
(10) Healthcare
Application
(11) Comput.
Cost
?Training and deploying LLMs are expensive as it requires considerable computing resources and memory…[…]? (arXiv, November 2023 (Zhang et al., 2023a))
(12) Benchmark
Contamination
Quantization
@@
?[…] Extensive experiments on DepEdit show that existing knowledge editing methods are sensitive to the surface form of knowledge, and that they have limited performance in inferring the implications of edited facts.? (arXiv, December 2023 (Li et al., 2023a))
While HDBSCAN provides an interpretable, unsupervised clustering approach, it has several limitations for trend analysis. HDBSCAN assigns each paper to only one cluster and becomes increasingly fragmented as more clusters are added (especially in ACL) making it difficult to capture overlapping limitations or analyze trends at finer topical granularity. To address these issues, we now turn to the LlooM approach, which supports multi-label assignment and allows for broader topical coverage. The following section presents the main findings using LlooM, including trend analysis of LLLMs topics; additional trend results from HDBSCAN are included in Section C in the supplementary material for completeness.
4.4. Clustering Results (LlooM)
4.4.1. Topics Identified within ACL and arXiv
Key Insights LlooM identifies a more fine-grained set of topics than HDBSCAN: 13 topics for ACL, including new topics like Multimodality, Language and Culture, and Knowledge Editing; and 15 for arXiv, adding concerns such as Alignment, Trustworthiness, and Generalization. Reasoning, Generalization, and Hallucination are top-ranking topics in both ACL and arXiv. Beyond these shared concerns, ACL places additional emphasis on Knowledge Editing, while arXiv is led by Trustworthiness and Alignment.
Figure 8 shows the distribution of LLM limitation topics in the ACL and arXiv datasets as identified by the LlooM method. The share of outliers is 7.5% in ACL and 6.7% in arXiv.
In ACL, Reasoning remains the most prominent topic (26.3%), followed by Generalization (24.8%) and Knowledge Editing (21.2%). Compared to HDBSCAN, the distribution across top limitations appears more balanced and can be explained by the fact that LlooM allows papers to belong to multiple clusters. ArXiv is led by Trustworthiness (21.0%), followed by Reasoning (13.2%), Generalization (10.1%), as well as Alignment Limitations, Hallucination, and Bias and Fairness, which appear at nearly equal levels ($\sim$7.9%).
A number of topics are specific to LlooM. These are listed below, with explanations based on the prompts used by LlooM to guide topic assignment (see Table 14 and Table 15 in the supplementary material). For each cluster, we also provide representative paper examples in Table 9.
  • Trustworthiness (arXiv): the largest cluster in the arXiv set (Figure 8). This category describes concerns about the reliability, transparency, and reproducibility of LLM outputs (see example 20 which refers to concerns about the reliability of outputs generated by LMs). As a broad category, it often overlaps with related issues like hallucination or alignment, as discussed further in Section D of the supplementary material.
  • Generalization (arXiv): Previously ACL-only in HDBSCAN, this cluster now appears in arXiv as well. It captures failures across domains or tasks; e.g. example 19 shows reliance on shallow heuristics over true generalization.
  • Language and Cultural Limitations (ACL): difficulties in handling multilingual input, low-resource languages, or culturally specific content, previously grouped under the Reasoning cluster in HDBSCAN.
  • Knowledge Editing (ACL & arXiv): efforts to modify, correct, or update model knowledge post-training: for instance, example 15 (ACL) in Table 9 introduces methods for editing factual content without retraining.
  • Privacy Risks (ACL & arXiv): previously part of the Security cluster in HDBSCAN for both datasets, this now is a distinct category in LlooM. It captures risks of leaking sensitive training data via model outputs or queries. Example 21 (arXiv) demonstrates privacy breaches from pretraining on sensitive data.
  • Multimodality (ACL): challenges in integrating and reasoning over inputs from different modalities, such as text and images (example 16, ACL), previously grouped under the Reasoning cluster in HDBSCAN.
  • Computational Cost (ACL): the high memory, compute, and energy demands of training or deploying LLMs (example 17).
In summary, the LlooM clustering identifies a broader set of limitation types than HDBSCAN, including implementation-level issues (e.g., Privacy, Computational Cost) and behavioral factors (e.g., Overconfidence, Prompt Sensitivity) that were previously grouped under broader categories. It also reflects some dataset-specific shifts, such as Generalization appearing in arXiv and Data Contamination in ACL. In the next section, we examine how attention to these topics has changed over time.
Table 9. Representative limitation-focused papers from ACL and arXiv, grouped by topic. Only topics identified by LlooM and absent from HDBSCAN-based analyses are included. The numbers in parentheses in the ?Topic? column indicate the example ID referenced in the main text.
Topic Evidence
(15) Knowledge
Editing
?Even the most advanced language models remain susceptible to errors necessitating to modify these models without initiating a comprehensive retraining process.? (EMNLP 2023 (Akyürek et al., 2023))
@@ 4.4.1. Topics Identified within ACL and arXiv
?However, their large size and computational demands, coupled with privacy concerns in data transmission, limit their use in resource-constrained and privacy-centric settings.? (NAACL 2024 (Jiang et al., 2024b))
(18) Data
Contamination
(19) Generalization ?We suggest that the lack of generalization […] means that the PLMs are currently not learning NLI, but rather spurious heuristics.? (arXiv, January 2022 (Gubelmann and Handschuh, 2022))
(20) Alignment
& Trustworthiness
(21) Privacy
Risks
?LLMs […] have been shown to memorize instances of training data thereby potentially revealing private information processed during pre-training.? (arXiv, May 2022 (Majmudar et al., 2022))

### v3

@@ 3.4. Models and Prompting Evaluation
Similar to human annotators highlighting evidence in abstracts to indicate discussions of LLLMs, models are prompted to extract supporting evidence from the text. Each prompt instructs the model: ?Please respond in the following format, providing a rating and supporting evidence for the discussion of LLM limitations in each abstract. Do not include explanations, only cite the evidence found in the abstract.? Prompt 3, included as the most comprehensive, is available in Figure 1 of the online supplementary material.
@@ 3.5. Clustering
To ensure meaningful clusters and minimize spurious outliers, we tune UMAP, HDBSCAN and BERTopic parameters separately for ACL and arXiv. This separation accounts for differences in corpus size and topical heterogeneity. We further apply a distance-based outlier reassignment strategy after observing that many outliers were not true noise but semantically close to existing clusters, suggesting misclassification by HDBSCAN. Full parameter settings and details of the reassignment procedure are provided in Section B of the online supplementary material.
@@
Clustering Validation
We assess the robustness of the identified limitation topics using (i) cross-method comparison, which evaluates topic-level alignment (Jaccard overlap, AMI) and trend-level alignment (Kendall’s Tau, Spearman’s $\rho$) between HDBSCAN and LlooM, and the validity of these topics using (ii) human evaluation, in which two expert annotators independently re-annotate papers with limitation topics predefined by HDBSCAN and LlooM, and agreement with each clustering method is measured, based on a stratified sample of 50 ACL and 60 arXiv papers covering all topics identified by both methods. Detailed procedures and results are reported in 4.4.
@@ 4.1. LLM-based filtering evaluation
(1) ?Unlocking Adversarial Suffix Optimization Without Affirmative Phrases: Efficient Black-box Jailbreaking via LLM as Optimizer? (Jiang et al., 2024),
@@ 4.2. LLM and LLLMs Trends Over Time
In both corpora, the (i) overall share of LLM-related papers has grown substantially since early 2023. This trend is particularly steep in ACL, where, by late 2024, over 75% of ACL papers are related to LLMs.
@@
$i$ Proportion of LLM limitation papers among all LLM papers.
@@
This suggests a notable shift in NLP research, with LLMs becoming central to the field. In arXiv, growth is more moderate but consistent, hitting just above 30% of papers by the end of the same period. The lower proportion in arXiv might be due to different levels of engagement with LLM research across the categories in our study, as shown in Figure 4 in the online supplementary material. In cs.CL, LLMs are widely discussed, reaching around 80% by early 2025, similar to ACL, while cs.AI shows a sharper but lower rise, peaking around around 50–60%. In contrast, in areas like cs.CV and cs.LG, their presence remains below 20%.
@@
In the following sections, we refine this analysis and examine these emerging discussions in detail through topic clustering. We discuss topics identified with HDBSCAN and LlooM, continue with trends identified with LlooM and provide trends based on HDBSCAN trends in the online supplementary material.
4.3. Clustering Results
4.3.1. Topics identified within ACL and arXiv with HDBSCAN and LlooM
Key Insights Both methods consistently identify shared high-level limitation categories, including Reasoning, Hallucination, Security Risks, Social Bias, Generalization, and Long-Context limitations. Differences between HDBSCAN and LlooM primarily reflect methodological choices: HDBSCAN concentrates papers into fewer, broader clusters, whereas LlooM’s multi-label assignments distribute papers across finer-grained limitation categories.
@@ 4.3.1. Topics identified within ACL and arXiv with HDBSCAN and LlooM
Figures 6 and 7 show topic distributions for ACL and arXiv under HDBSCAN+BERTopic and LlooM, respectively. HDBSCAN identifies 7 topics in ACL and 15 in arXiv (13.1% and 14.1% outliers), while LlooM produces comparable high-level topics with lower outlier rates (7.5% and 6.7%).
Both clustering methods identify Reasoning as the dominant limitation topic in ACL (36.4% under HDBSCAN vs. 26.3% under LlooM). HDBSCAN concentrates a larger share of papers in this single topic, while LlooM distributes papers more evenly across top limitations, including Generalization and Knowledge Editing, due to its multi-label assignment. In arXiv, both approaches show a broader topical spread than in ACL, with the most frequent themes being Social Bias (16.8%), Security Risks (15.9%), and Reasoning (10.9%) under HDBSCAN, and Trustworthiness (21.0%), Reasoning (13.2%) and Generalization (10.1%) under LlooM.
Topics shared across both methods
Table 7 presents representative papers for topics shared between HDBSCAN and LlooM in the ACL and arXiv datasets. Tables 6 and 7 in the online supplementary material report the dominant terms (top 20 keywords) for each limitation cluster identified with HDBSCAN.
Under both clustering methods, the Reasoning cluster captures core cognitive tasks such as natural language understanding (NLU), inference, and logical problem-solving, spanning a broad range of reasoning types (e.g.  temporal and commonsense). It also includes prompt-related terms such as chain-of-thought and prompt, highlighting the role of prompt engineering in complex reasoning. Under HDBSCAN in ACL, however, the Reasoning cluster is broader, absorbing multimodal tasks (example 1 in Table 7), and overlapping with multilinguality and benchmark design. In contrast, in arXiv under HDBSCAN and in both datasets under LlooM, Reasoning is more distinct, as Multilinguality / Language and Cultural Limitations (difficulties in handling multilingual input, low-resource languages, or culturally specific content) and Multimodality (challenges in integrating and reasoning over inputs from different modalities, such as text and images, example 9) are separated into their own clusters. This difference likely contributes to the higher prevalence of Reasoning in ACL under HDBSCAN, though it may also partly reflect clustering artefacts.
@@
Several shared clusters focus on the correctness and reliability of model outputs in both ACL and arXiv. The Hallucination cluster addresses failures in factual accuracy, with terms related to faithfulness, trust, and correctness (example 2), and also appears in multimodal contexts such as image captioning and generation. The Security cluster captures risks including jailbreaks, adversarial prompting, and backdoors (example 3). Under LlooM, security-related concerns further separate into distinct Security and Privacy topics. The Social Bias cluster focuses on fairness, representation, and demographic disparity (example 5), with keywords such as stereotype, gender, cultural, and demographic.
@@
Other shared clusters reflect technical constraints in model design and deployment. Long Context / Context & Memory Limitations address failures in handling extended inputs and maintaining information over long outputs (example 6 in Table 7), closely related to increased resource demands captured by the Computational Cost cluster (example 10). The Generalization cluster captures limitations in robustness and domain transfer (example 4). Knowledge Editing focuses on the difficulty of updating or correcting model knowledge without retraining, often leading to unintended performance degradation (example 8).
Table 7. Representative limitation-focused papers for topics that are shared by HDBSCAN and LlooM and that occur in the ACL and/or arXiv corpora. Topics that occur in both corpora are shown once, with dataset indicated by the venue/date preceding the citation. Unless marked with $\dagger$, each paper is assigned to the conceptually corresponding topic by both methods. Topic labels follow HDBSCAN; parentheses show the LlooM label when it differs. Numbers in the ?Topic? column indicate the example ID referenced in the main text.
@@
(2) Hallucination ?Challenges on hallucination and factual inconsistency continue to impede their [LLMs’] wider real-world adoption. […] However, challenges remain, particularly regarding… generating information not present in the evidence (hallucination).? (EACL 2024 (Lv et al., 2024))
Security Risks
?The prevalence and strong capability of LLMs present significant safety and ethical risks if exploited by malicious users. […] Experiments reveal that our attacks effectively compromise the performance of all detectors in the study with plausible generations […].? (TACL 2024 (Shi et al., 2024))
@@
?However, in-domain demonstrations are not always readily available in real scenarios, leading to cross-domain in-context learning. Besides, LLMs are still facing challenges in long-tail knowledge in unseen and unfamiliar domains.? (EMNLP 2023 (Long et al., 2023))
(5) Social Bias
(Bias & Fairness)
@@
(7) Multiling.
(Lang. & Cult.)
?The automated systems […] are primarily designed for and work far more effectively in English than in the world’s other 7,000 languages. […]? (arXiv, June 2023 (Nicholas and Bhatia, 2023))
@@
?Existing editing methods lead to inevitable performance deterioration on general benchmarks. […] When the number of edits is slightly large, the intrinsic knowledge structure of the model is disrupted or even completely damaged.? (arXiv, October 2024 (Li et al., 2024a))
@@
?Training and deploying LLMs are expensive as it requires considerable computing resources and memory…[…]? (arXiv, November 2023 (Zhang et al., 2023b))
$\dagger$ Assigned to Reasoning under LlooM.
HDBSCAN-specific clusters
In addition to shared topics, HDBSCAN identifies several clusters that do not emerge under LlooM. Representative papers for these clusters are shown in Table 8 in the online supplementary material. In ACL, the Uncertainty cluster describes behavioral instability, including prompt sensitivity and calibration errors. In arXiv, additional clusters capture a wider range of specialized concerns: Conversational Limitations Code Generation, Healthcare Application, Benchmark Contamination, and Quantization. These clusters reflect both domain-specific challenges (e.g. Healthcare Application, which focuses on the use of LLMs in clinical settings and is primarily concerned with their black-box nature, see example 13) and implementation-level trade-offs. For example, Benchmark Contamination refers to test data leakage into training sets (example 14), and Quantization concerns scaling efficiency (example 15).
LlooM-specific clusters
A number of topics are specific to LlooM. These are listed below, with explanations based on the prompts used by LlooM to guide topic assignment (see Table 3 and Table 4 in the online supplementary material). For each LlooM-specific cluster, we also provide representative paper examples in Table 8 in the online supplementary material.
  • Trustworthiness (arXiv): the largest cluster in the arXiv set (Figure 7). This category describes concerns about the reliability, transparency, and reproducibility of LLM outputs (see example 19 and 20 which refer to concerns about the reliability of outputs generated by LMs). As a broad category, it often overlaps with related issues like hallucination or alignment, as discussed further in Section D of the online supplementary material.
  • Privacy Risks (ACL & arXiv): previously part of the Security cluster in HDBSCAN for both datasets, this now is a distinct category in LlooM. It captures risks of leaking sensitive training data via model outputs or queries. Example 16 (ACL) demonstrates privacy breaches from pretraining on sensitive data.
The comparison of shared and method-specific clusters highlights both similarities and systematic differences between HDBSCAN and LlooM (we quantify these differences in 4.4). While HDBSCAN provides an interpretable, unsupervised clustering of limitations, it assigns each paper to a single cluster and becomes increasingly fragmented as more clusters are added, particularly in the ACL dataset. This makes it difficult to capture overlapping limitations and may complicate the interpretation of trends. By contrast, LlooM allows for broader, non-mutually exclusive categories through multi-label assignment, which reduces fragmentation and provides a more stable basis for longitudinal analysis. Accordingly, we rely on LlooM for the trend analysis in the following section. Trend results derived from HDBSCAN are reported in Section C of the online supplementary material for completeness.